A 2.80 kg block (mass 1) and a 4.90 kg block (mass 2) are connected by a light string as shown; the inclination of the ramp is 40.0°. Friction is negligible.

40°

(a) What is the acceleration (in m/s2) of each block? (Enter the magnitude.)

_____ m/s2

(b) What is the tension (in N) in the string?

_____ N
Answer #1

Tension in the string T = m2 g =(4.9)(9.8) = 48.02 N

Net force on the system F = T -(m1g sin 40)

= 48.02 -(2.8x9.8xsin40)

= 48.02-17.638 = 30.3819 N

Acceleration of the system a = F /(m1+m2)

= 30.3819 /(2.8+4.9)

= 3.945 m/s 2